# RLEF: Grounding Code LLMs in Execution Feedback with Reinforcement Learning

### v1

@@ Abstract
Large language models (LLMs) deployed as agents solve user-specified tasks over multiple steps while keeping the required manual engagement to a minimum. Crucially, such LLMs need to ground their generations in any feedback obtained to reliably achieve desired outcomes. We propose an end-to-end reinforcement learning method for teaching models to leverage execution feedback in the realm of code synthesis, where state-of-the-art LLMs struggle to improve code iteratively compared to independent sampling. We benchmark on competitive programming tasks, where we achieve new start-of-the art results with both small (8B parameters) and large (70B) models while reducing the amount of samples required by an order of magnitude. Our analysis of inference-time behavior demonstrates that our method produces LLMs that effectively leverage automatic feedback over multiple steps.
@@ 1 Introduction
We posit that any decision-making agent offering a natural language interface has to possess two skills. First, the ability to accurately deduce a user’s intent when prompted; for LLMs, this is typically achieved by fine-tuning to follow instructions according to user preferences (Ouyang et al., 2022; Rafailov et al., 2023). Second, feedback on intermediate results of the agent’s actions has to be taken into account to arrive at the desired outcome. For example, a web page containing a necessary bit of information might have gone offline, requiring another search engine query. In the context of code generation, feedback can provide information about implementation bugs as well as constraints that are inefficient or cumbersome to specify in full detail, e.g., software and hardware platform details or library dependencies. Intermediate feedback is therefore crucial to ground LLM generations in the concrete situations encountered at inference time.
@@
Figure 2: Left: Overview of reinforcement learning with execution feedback (RLEF). The LLM is repeatedly prompted to implement code according to a problem description. Each attempt is evaluated on a public test set; upon failure, feedback is inserted into the conversation. If public tests are passing, or a specified turn limit is reached, execution on additional, private tests determines the reward signal. The model is then updated to optimize the reward with PPO. Right: Example dialog with two model responses. Execution feedback hints at an inefficient first solution, to which the model responds to utilizing a cache. The code passing the public test sets will be evaluated on the full test set.
@@
In this work, we aim to endow pre-trained LLMs with the aforementioned skills, task alignment and grounding in inference-time feedback, in the domain of code synthesis from natural language descriptions (Chen et al., 2021; Rozière et al., 2023). Here, feedback is naturally provided as the result of the execution of generated code in the form of error messages and unit test results. However, to date, utilizing such feedback for code generation with LLMs has failed to yield substantial improvements when taking computational demands into account; indeed, obtaining samples independently often results in higher accuracy for a fixed inference budget (Kapoor et al., 2024; Xia et al., 2024). As a test-bed to investigate and improve grounding in execution feedback, we propose to frame code generation as an iterative task, repeatedly asking an LLM to produce code according to a provided natural language description (Fig. 2). After each generation, code is evaluated on example test cases and the resulting feedback is provided as additional context for subsequent attempts. We thus obtain an interactive environment where actions correspond to code and observations correspond to execution feedback. Importantly, such a framing permits end-to-end optimization with reinforcement learning (RL) algorithms to maximize a reward signal – here, a binary reward based on whether the final code solution passes a set of held-out test cases.
@@ 2.1 Iterative Code Synthesis
Our conversation flow for code generation is depicted in Fig. 2. Concretely, we start the dialog with the problem description and query the LLM for an initial solution. The solution is verified against the public test set, which yields results in the form of passed and failed test cases, as well as potential syntax or runtime errors. If any public test fails, this execution feedback is formatted and appended to the dialog The LLM is then queried for an updated code solution, with the original problem text, previous solutions and their respective feedback provided in the prompt. If the solution passes all public tests, or a specified turn limit is reached, it is considered to be final and will be submitted for evaluation on the private test set. The kind reader is referred to Appendix C for a listing of our prompt and execution feedback templates.
@@ 2.2 Reinforcement Learning with Execution Feedback
For optimizing a policy in the above environment we employ Proximal Policy Optimization (PPO), a common choice for fine-tuning large language models (Schulman et al., 2017; Ziegler et al., 2020; Ouyang et al., 2022). Following previous work, we include a KL penalty in our reward signal, acting both as an entropy bonus and as regularization towards the distribution of the LLMs we start from. In initial experiments we found that a possible failure mode concerns the generation of invalid code in non-final responses, which we address by providing a small penalty for invalid responses. Denoting the policy to be optimized with $\pi$ and the initial policy with $\rho$, and abbreviating previous observations and actions with $c_{t}=o_{0},a_{0},o_{1},a_{1},\dots,o_{t}$ our reward function at step $t$ is thus
@@
with a constant $\beta$ trading off between task reward and KL maximization. For PPO, we compute policy gradients by incorporating a concurrently learned value function as a baseline, i.e., we train the policy to maximize the advantage $A_{t}=R(s_{t},a_{t})-V(c_{t})$.
@@ 3.1 Setup
We perform experiments on the CodeContests benchmark introduced by Li et al. (2022) which requires generating a code solution to a problem specified in natural language along with a textual description of public test cases. Problems are of high difficulty and used in human competitive programming with a focus on algorithms, data structures and runtime efficiency. The correctness of solutions is evaluated with private tests that are hidden from contestants, which we implement in our setup by presenting feedback from public tests only. The CodeContests dataset consists of a training set and two test sets, “valid” and “test”, consisting of 117 and 165 problems, respectively; we use the former for model and hyperparameter selection. We optimize our models on the training set, from which we discard 115 of the 13,328 problems due to missing public test cases. We prompt and train all models to output Python 3 code.
@@ 3.2 Main Results
AlphaCode 9B Li et al. (2022) $10@1k$ $16.9$ $13.3$
AlphaCode 41B + clustering Li et al. (2022) $10@1k$ $21.0$ $16.4$
Code Llama 34B + PPO Xu et al. (2024) $10@1k$ $19.7$ $22.4$
@@ 3.3 Inference-time Behavior
In Section 3.3 we first take a closer look at single- and multi-turn performance with a fixed budget of 3 LLM generations (1@3). This corresponds to our iterative setup with up to three model responses, or three independent responses for single-turn results. We further consider generalization to two popular code generation benchmarks, HumanEval+ and MBPP+ (Liu et al., 2023), which we modify to match our iterative code generation setup with “base” tests for inference-time execution feedback and “plus” tests for solve rate estimation (see Section C.4 for details). Our results demonstrate that, when considering a fixed sample budget, base models rarely benefit from access to faulty solutions and execution feedback in the multi-turn code generation setup. This also applies to gpt-4o-2024-05-13, which shows stronger performance when sampling solutions independently on CodeContests and HumanEval+. After RLEF training, the 8B and 70B Llama 3.1 model both benefit from execution feedback and can therefore achieve larger gains on top of improved single-turn scores, with the exception of the 8B model on CodeContests and MBPP+ where single-turn performance drops. While multi-turn gains from RLEF are most pronounced on CodeContests, the training domain of our models, we also observe notable improvements on HumanEval+ and MBPP+.
@@
Figure 4: (a) Pass@1 and pass@10 across turn limits with RLEF-trained models, providing either true or random execution feedback (temperature 0.2). With random feedback pass@1 is reduced while pass@10 suffers only slightly, indicating that programs can be repaired less consistently. (b) Impact of turn limits on 10@k solve rates per sample budget (top: 8B model, bottom: 70B model) with temperature 1.0. With RLEF, iterative code generation can leverage up to 5 turns to achieve best performance with 5 turns across.
@@ 3.4.2 Single-turn Training
In Section 3.4 we compare our iterative code generation setup to traditional, single-turn generation where the model is not presented with inference-time feedback. We use the same training loop for single generations, albeit without the penalty for invalid code (Section 2.2) as this is subsumed by the reward signal for incorrect solutions. For Llama 3.1 Instruct 8B, single-turn training (ST) hurts performance on the test set. The 70B model benefits from single-turn training and improves over multi-turn SFT results in Section 3.4. Moreover, we observe transfer in that applying the model trained for single turns in a multi-turn setting improves 1@3 solve rates. We attribute this to the existent but comparabily weak multi-turn capabilities of the vanilla 70B Instruct model. Overall, we see strongest performance with the RLEF method employing multiple turns at training and inference time.
@@ 4 Related Work
More recently, several works investigated prompting and flow engineering techniques to improve performance at inference time, including the verification of generated code via compilation and execution, followed by re-prompting. Shinn et al. (2023) and Chen et al. (2024) use feedback from unit tests to correct previously wrong generations and found it crucial to include model-generated error analysis in the prompt for successive generations. LDB (Zhong et al., 2024), AlphaCodium (Ridnik et al., 2024) and MapCoder (Islam et al., 2024) can be regarded as agentic frameworks as they provide rich manual scaffolding for code generation, chaining several LLM calls (e.g., for chain-of-thought planning, test generation, and program repair) combined with code execution. These approaches are effective on difficult benchmarks, such as the CodeContests dataset we consider in this work, but significantly increase inference cost by requiring dozens of LLM calls per solution.
@@
Fine-tuning large language models with reinforcement learning is a popular method for aligning their output to user preferences (Ziegler et al., 2020; Touvron et al., 2023; OpenAI, 2023; DeepSeek-AI et al., 2024; AI @ Meta, 2024). Here, the learning signal is provided by special-purpose reward models. For code synthesis, however, rewards can be determined by executing LLM generations against available test cases. Le et al. (2022) pre-train an LLM for code generation and subsequently fine-tune it with both policy gradients and next-token loss on rewards from execution. For policy rollouts, they perform program refinement and repair to increase the likelihood of sampling correct programs; however, their model is trained on final solutions only. Shojaee et al. (2023) utilize a rich reward signal that considers execution as well as similarity to ground truth code, introducing a dependency on human-provided solutions. Finally, Xu et al. (2024) fine-tune a stronger, code-specific LLM in a simpler setup with a binary reward from unit tests and observe substantial improvements from RL on the difficult competitive programming benchmark we consider here. We likewise propose a simple setting without extra inference scaffolding or usage of ground truth solutions. Crucially, we expand the traditional natural-language-to-code setting to an iterative environment where execution feedback is not only provided as a scalar reward but also in textual form. This allows us to shift focus from large-sample inference regimes to obtaining high accuracy with low sample budgets. Concurrently to our work, Kumar et al. (2024) propose a two-stage RL method (SCoRe) to improve the self-correction capabilities of LLMs and train them to output two successive solutions. In contrast to our method, SCoRe does not leverage execution feedback at inference time and instead asks the model to reconsider its initial solution. While this approach allows for potential applications to domains where automatic feedback is not available, it cannot benefit from the information provided in the feedback message. Furthermore, inference-time feedback can help the model generalize to new environments after training.
@@ 5 Conclusion
While our results demonstrate effective usage of inference-time feedback, the code synthesis task we consider is limited to improving a single solution to a given problem. Generalizing our method to environments with larger tasks that require decomposition, either via manual scaffolding or, eventually, in a self-directed manner, remains the subject of further research.
@@ Acknowledgements.
We thank Quentin Carbonneaux, Chris Cummins, Olivier Duchenne, Fabian Gloeckle, Baptiste Roziere, Sten Sootla, Nicolas Usunier, and Sida Wang for helpful technical contributions, suggestions, and insightful discussions.
@@ A.1 RLEF
We initialize separate policy and value function networks from pre-trained and instruction-tuned LLMs as indicated in the respective experiments; for the value function, we replace the output layer with a randomly initialized linear projection. For PPO, we use a AdamW (Loshchilov & Hutter, 2019) with a learning rate of $2e^{-7}$, weight decay of 0.1, and a linear warm-up over 50 steps. We set the KL regularization factor $\beta$ of the reward term to $0.05$ (Section 2.2). All models are trained with an online, asynchronous training infrastructure that decouples inference and optimization. We incorporate importance sampling in PPO’s clipped surrogate objective (Schulman et al., 2017, Eq.7):
@@
$L^{CLIP}(\theta)$ $=\mathbb{\hat{E}}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\operatorname{clip}\left(r_{t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{t}\right)\right]$
for model parameters $\theta$, advantage $\hat{A}_{t}$, and the behavior policy $\pi_{\text{b}}$. We set $\epsilon=0.2$. For optimizing the value function, we use a clipped value loss.

### v2

@@ Abstract
Large language models (LLMs) deployed as agents solve user-specified tasks over multiple steps while keeping the required manual engagement to a minimum. Crucially, such LLMs need to ground their generations in any feedback obtained to reliably achieve the desired outcomes. We propose an end-to-end reinforcement learning method for teaching models to leverage execution feedback in the realm of code synthesis, where state-of-the-art LLMs struggle to improve code iteratively compared to independent sampling. We benchmark on competitive programming tasks, where we achieve new state-of-the-art results with both small (8B parameters) and large (70B) models while reducing the amount of samples required by an order of magnitude. Our analysis of inference-time behavior demonstrates that our method produces LLMs that effectively leverage automatic feedback over multiple steps.
@@ 1 Introduction
We posit that any decision-making agent offering a natural language interface has to possess two skills. First, the ability to accurately deduce a user’s intent when prompted; for LLMs, this is typically achieved by fine-tuning to follow instructions according to user preferences (Ouyang et al., 2022; Rafailov et al., 2023). Second, feedback on intermediate results of the agent’s actions has to be taken into account to arrive at the desired outcome. For example, a web page that contains a necessary bit of information might have gone offline, requiring another search engine query. In the context of code generation, feedback can provide information about implementation bugs as well as constraints that are inefficient or cumbersome to specify in full detail, e.g., software and hardware platform details or library dependencies. Intermediate feedback is therefore crucial to ground LLM generations in the concrete situations encountered at inference time.
@@
Figure 2: Left: Overview of reinforcement learning with execution feedback (RLEF). The LLM is repeatedly prompted to implement code according to a problem description. Each attempt is evaluated on a public test set; upon failure, feedback is inserted into the conversation. If public tests are passing, or a specified turn limit is reached, execution on additional, private tests determines the reward. The model is then updated with PPO. Right: Example dialog with two model responses. Execution feedback hints at an inefficient first solution, to which the model responds to utilizing a cache. The code passing the public test sets will be evaluated on the full test set.
@@
In this work, we aim to endow pre-trained LLMs with the aforementioned skills, task alignment and grounding in inference-time feedback, in the domain of code synthesis from natural language descriptions (Chen et al., 2021; Rozière et al., 2023). Here, feedback is naturally provided as the result of the execution of generated code in the form of error messages and unit test results. However, to date, utilizing such feedback for code generation with LLMs has failed to yield substantial improvements when taking computational demands into account; indeed, obtaining samples independently often results in higher accuracy for a fixed inference budget (Kapoor et al., 2024; Xia et al., 2024). As a test bed to investigate and improve grounding in execution feedback, we propose to frame code generation as an iterative task, repeatedly asking an LLM to produce code according to a provided natural language description (Fig. 2). After each generation, code is evaluated on example test cases and the resulting feedback is provided as additional context for subsequent attempts. We thus obtain an interactive environment where actions correspond to code and observations correspond to execution feedback. Importantly, such a framing permits end-to-end optimization with reinforcement learning (RL) algorithms to maximize a reward signal – here, a binary reward based on whether the final code solution passes a set of held-out test cases.
@@ 2.1 Iterative Code Synthesis
Our conversation flow for code generation is depicted in Fig. 2. Concretely, we start the dialog with the problem description and query the LLM for an initial solution. The solution is verified against the public test set, which yields results in the form of passed and failed test cases, as well as potential syntax or runtime errors. If any public test fails, this execution feedback is formatted and appended to the dialog. The LLM is then queried for an updated code solution, with the original problem text, previous solutions, and their respective feedback provided in the prompt. If the solution passes all public tests, or a specified turn limit is reached, it is considered to be final and will be submitted for evaluation on the private test set. The kind reader is referred to Appendix C for a listing of our prompt and execution feedback templates.
@@ 2.2 Reinforcement Learning with Execution Feedback
For optimizing a policy in the above environment we employ Proximal Policy Optimization (PPO), a common choice for fine-tuning large language models (Schulman et al., 2017; Ziegler et al., 2020; Ouyang et al., 2022). Following previous work, we include a KL penalty in our reward signal, acting both as an entropy bonus and as regularization towards the distribution of the LLMs we start from. In initial experiments we found that a possible failure mode concerns the generation of invalid code in non-final responses, which we address by providing a small penalty for invalid responses. Denoting the policy to be optimized with $\pi$ and the initial policy with $\rho$, and abbreviating previous observations and actions with $c_{t}=o_{0},a_{0},o_{1},a_{1},\dots,o_{t}$ our reward function at step $t$ is
@@
with a constant $\beta$ trading off between task reward and KL maximization. For PPO, we compute policy gradients by incorporating a concurrently learned value function as a baseline, i.e., we train the policy to maximize the advantage $A_{t}=-V(c_{t})+\sum_{i=t}^{T}R(s_{i},a_{i})$; see Section A.1.
@@ 3.1 Setup
We perform experiments on the CodeContests benchmark introduced by Li et al. (2022) which requires generating a code solution to a problem specified in natural language along with a textual description of public test cases. Problems are of high difficulty and used in human competitive programming with a focus on algorithms, data structures and runtime efficiency. The correctness of solutions is evaluated with private tests that are hidden from contestants, which we implement in our setup by presenting feedback from public tests only. The CodeContests dataset consists of a training set and two test sets, “valid” and “test”, consisting of 117 and 165 problems, respectively; we use the former for model and hyperparameter selection. We optimize our models on the training set, from which we discard 669 of the 13,328 problems due to missing public or private test cases. We prompt and train all models to output Python 3 code.
@@ 3.2 Main Results
AlphaCode 9B Li et al. (2022) $10$@$1000$ $16.9$ $13.3$
AlphaCode 41B + clustering Li et al. (2022) $10$@$1000$ $21.0$ $16.4$
Code Llama 34B + PPO Xu et al. (2024) $10$@$1000$ $19.7$ $22.4$
@@ 3.3 Inference-time Behavior
In Section 3.3 we first take a closer look at single- and multi-turn performance with a fixed budget of 3 LLM generations ($1$@$3$). This corresponds to our iterative setup with up to three model responses, or three independent responses for single-turn results. We further consider generalization to two popular code generation benchmarks, HumanEval+ and MBPP+ (Liu et al., 2023b), which we modify to match our iterative code generation setup with “base” tests for inference-time execution feedback and “plus” tests for solve rate estimation (see Section C.4 for details). Our results demonstrate that, when considering a fixed sample budget, base models rarely benefit from access to faulty solutions and execution feedback in the multi-turn code generation setup. This also applies to gpt-4o-2024-05-13, which shows stronger performance when sampling solutions independently on CodeContests and HumanEval+. After RLEF training, the 8B and 70B Llama 3.1 model both benefit from execution feedback and can therefore achieve larger gains on top of improved single-turn scores, with the exception of the 8B model on CodeContests and MBPP+ where single-turn performance drops. While multi-turn gains from RLEF are most pronounced on CodeContests, the training domain of our models, we also observe notable improvements on HumanEval+ and MBPP+.
@@
Figure 4: (a) Pass@1 and pass@10 across turn limits with RLEF-trained models, providing either true or random execution feedback (temperature 0.2). With random feedback pass@1 is reduced while pass@10 suffers only slightly, indicating that programs can be repaired less consistently. (b) Impact of turn limits on 10@k solve rates per sample budget (top: 8B model, bottom: 70B model) with temperature 1.0. With RLEF, iterative code generation can leverage up to 5 turns to achieve compute-optimal performance.
@@ 3.4.2 Single-turn Training
In Section 3.4 we compare our iterative code generation setup to traditional, single-turn generation where the model is not presented with inference-time feedback. We use the same training loop for single generations, albeit without the penalty for invalid code (Section 2.2) as this is subsumed by the reward signal for incorrect solutions. For Llama 3.1 Instruct 8B, single-turn training (ST) hurts performance on the test set. The 70B model benefits from single-turn training and improves over multi-turn SFT results in Section 3.4. Moreover, we observe transfer in that applying the single-turn model in a multi-turn setting improves $1$@$3$ solve rates. We attribute this to the existent but comparabily weak multi-turn capabilities of the vanilla 70B Instruct model. Overall, we see strongest performance with the RLEF method employing multiple turns at training and inference time.
Further ablations can be found in the appendix. In Section B.3 we evaluate the effect of training a dedicated repair model on outputs of the single-turn 8B training run in Section 3.4, similar to (Le et al., 2022). Together, the single-turn and repair model obtain $1$@$3$ solve rates of 14.8 on the validation set and 12.6 on the test set; an improvement over the single-turn model alone (10.2 and 10.9) but significantly below the corresponding multi-turn model (17.2 and 16.0). In Section B.4 we show that withholding public test execution feedback during training results in significantly worse performance. Finally, in Section B.5 we experimentally validate the design choice of a turn-level value function (Section 2.2).
@@ 4 Related Work
More recently, several works investigated prompting and flow engineering techniques to improve performance at inference time, including the verification of generated code via compilation and execution, followed by re-prompting. Shinn et al. (2023) and Chen et al. (2024b) use feedback from unit tests to correct previously wrong generations and found it crucial to include model-generated error analysis in the prompt for successive generations. LDB (Zhong et al., 2024), AlphaCodium (Ridnik et al., 2024) and MapCoder (Islam et al., 2024) can be regarded as agentic frameworks as they provide rich manual scaffolding for code generation, chaining several LLM calls (e.g., for chain-of-thought planning, test generation, and program repair) combined with code execution. These approaches are effective on difficult benchmarks, such as the CodeContests dataset we consider in this work, but significantly increase inference cost by requiring dozens of LLM calls per solution.
@@
Fine-tuning large language models with reinforcement learning is a popular method for aligning their output to user preferences (Ziegler et al., 2020; Touvron et al., 2023; OpenAI, 2023; DeepSeek-AI et al., 2024; AI @ Meta, 2024). Here, the learning signal is provided by special-purpose reward models. For code synthesis, however, rewards can be determined by executing LLM generations against available test cases (Le et al., 2022; Shojaee et al., 2023; Dou et al., 2024; Yu et al., 2024). Le et al. (2022) pre-train an LLM for code generation and subsequently fine-tune it with both policy gradients and next-token loss on rewards from execution. From the rollouts obtained during fine-tuning, they train further models for predicting test outcome labels and for mapping incorrect to ground truth solutions, which allows for inference-time code correction based on test results (“critic sampling”), albeit without explicitly presenting the output from execution. Subsequently, Liu et al. (2023a) extends this work with an extended, fine-grained reward function. Finally, Xu et al. (2024) fine-tune a stronger, code-specific LLM in a simpler setup with a binary reward from unit tests and observe substantial improvements from RL on the difficult competitive programming benchmark we consider here. We likewise propose a simple setting without extra inference scaffolding or usage of ground truth solutions. Crucially, we expand the natural-language-to-code setting to an iterative environment where execution feedback is not only provided as a scalar reward but also in textual form. This allows us to acquire both code synthesis and code repair capabilties with a single model, and to shift focus from large-sample inference regimes to obtaining high accuracy with low sample budgets.
Concurrently to our work, Kumar et al. (2024) propose a two-stage RL method (SCoRe) to improve the self-correction capabilities of LLMs and train them to output two successive solutions. In contrast to our method, SCoRe does not leverage execution feedback at inference time and instead asks the model to reconsider its initial solution. While this approach allows for potential applications to domains where automatic feedback is not available, it cannot benefit from the information provided in the feedback message. Furthermore, inference-time feedback can help the model generalize to new environments after training. Finally, Chen et al. (2024a) address code generation with human feedback and develop an appropriate supervised fine-tuning strategy based on training a separate code repair model. In our work, we effectively leverage automatically generated feedback, formatted in natural language, with a single model only.
@@ 5 Conclusion
While our results demonstrate effective usage of inference-time feedback, the code synthesis task we consider is limited to improving a single solution to a given problem. Generalizing our method to environments with larger tasks that require decomposition, either via manual scaffolding or, eventually, in a self-directed manner, remains the subject of further research. Iterating on the execution results of unit tests naturally requires test cases, which may not be readily available. We regard a potential combination with automatic unit test generation (Watson et al., 2020; Jain et al., 2024) as an interesting avenue for further experiments.
@@ Acknowledgements.
We thank Chris Cummins, Olivier Duchenne, Fabian Gloeckle, Baptiste Roziere, Sten Sootla, Nicolas Usunier, and Sida Wang for helpful technical contributions, suggestions, and insightful discussions.
@@ A.1 RLEF
We initialize separate policy and value function networks from pre-trained and instruction-tuned LLMs as indicated in the respective experiments; for the value function, we replace the output layer with a randomly initialized linear projection. For PPO, we use AdamW (Loshchilov & Hutter, 2019) with a learning rate of $2e^{-7}$, weight decay of 0.1, and a linear warm-up over 50 steps. We set the KL regularization factor $\beta$ of the reward term to $0.05$ (Section 2.2). All models are trained with an online, asynchronous training infrastructure that decouples inference and optimization. We incorporate importance sampling in PPO’s clipped surrogate objective (Schulman et al., 2017, Eq.7):
@@
$L^{\pi}(\theta)$ $=\mathbb{\hat{E}}_{t}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\operatorname{clip}\left(r_{t}(\theta),1-\epsilon,1+\epsilon\right)\hat{A}_{t}\right)\right]$
for model parameters $\theta$, normalized advantage $\hat{A}_{t}$, and the behavior policy $\pi_{\text{b}}$. We set $\epsilon=0.2$.
For optimizing the value function, we use a clipped value loss. With value model parameters $\psi$ and reward function $R(s_{t},a_{t})$ (see Section 2.2) we have
$R_{t}$ $=\sum_{i=t}^{T}\gamma^{i-t}R(s_{i},a_{i})$
$L^{V}(\psi)$ $=\mathbb{\hat{E}}_{t}\left[\frac{1}{2}\max\left(\left(V_{\psi}(c_{t})-R_{t}\right)^{2},\left(\operatorname{clip}\left(V_{\psi}(c_{t}),V_{\psi_{\text{old}}}(c_{t})-\alpha,V_{\psi_{\text{old}}}(c_{t})+\alpha\right)-R_{t}\right)^{2}\right)\right]$
where we set the discount factor $\gamma$ to 1 and the value clipping threshold $\alpha$ to 0.2.